# Is Value Learning Really the Main Bottleneck in Offline RL?

### v1

@@ 1 Introduction
Our main contribution is an analysis of the bottlenecks in offline RL as evaluated via data-scaling properties of various algorithmic choices. Contrary to the conventional belief that value learning is the bottleneck of offline RL algorithms, we find that the performance is often limited by the choice of a policy extraction objective and the degree to which the policy generalizes at test time. This suggests that, with an appropriate policy extraction procedure (e.g., gradient-based policy extraction) and an appropriate recipe for handling policy generalization (e.g., test-time training with the value function), collecting more high-coverage data to train a value function is a universally better recipe for improving offline RL performance, whenever the practitioner has access to collecting some new data for learning. These results also imply that more research should be pursued in developing policy learning and generalization recipes to translate value learning advances into performant policies.
@@ 3 Main hypothesis
Our primary goal is to understand when and how the performance of offline RL can be bottlenecked in practice. As discussed earlier, there exist three potential factors that could bottleneck an offline RL algorithm: (B1) imperfect value function estimation from data, (B2) imperfect policy extraction from the learned value function, and (B3) imperfect policy generalization on the test-time states that the policy visits in evaluation rollouts. We note that the bottleneck of an offline RL algorithm under a certain dataset can always be attributed to one or some of these factors, since the policy will attain optimal performance if both value learning and policy extraction are perfect, and perfect generalization to test-time states is possible.
Our main hypothesis in this work is that, somewhat contrary to the prior belief that the accuracy of the value function is the primary factor limiting performance of offline RL methods, policy learning is often the main bottleneck of offline RL. In other words, while value function accuracy is certainly important, how the policy is extracted from the value function (B2) and how well the policy generalizes on states that it visits at the deployment time (B3) are often the main factors that significantly affect both the performance and scalability of offline RL. To verify this hypothesis, we conduct two main analyses in this paper. In Section 5, we compare the effects of value learning and policy extraction on performance under various types of environments, datasets, and algorithms (B1 and B2). In Section 6, we analyze the degree to which the policy generalizes on test-time states affects performance (B3).
@@ 5.2 Results: Policy extraction mechanisms substantially affect data-scaling trends
Figure 1 shows the data-scaling matrices of three policy extraction algorithms (AWR, DDPG+BC, and SfBC) and three value learning algorithms (IQL and {SARSA or CRL}) on eight environments, aggregated from a total of $7744$ runs ($4$ seeds for each cell). In each matrix, we individually tune the hyperparameter for policy extraction ($\alpha$ or $N$) for each entry. These matrices show how performance varies with different amounts of data for the value and the policy. In our analysis, we specifically focus on the color gradients of these matrices, which reveal the main limiting factor behind the performance of offline RL in each setting. Note that the color gradients are mostly either vertical, horizontal, or diagonal. Vertical ( ) color gradients indicate that the performance is most strongly affected by the amount of policy data, horizontal ( ) gradients indicate it is mostly affected by value data, and diagonal ( ) gradients indicate both.
@@
Among policy extraction algorithms, we find that DDPG+BC almost always achieves the best performance and scaling behaviors across the board, followed by SfBC, and the performance of AWR falls significantly behind the other two extraction algorithms in many cases. Notably, the data-scaling matrices of AWR always have vertical ( ) or diagonal ( ) color gradients, implicitly implying that it does not fully utilize the value function (see Section 5.3 for clearer evidence). In other words, a non-careful choice of the policy extraction algorithm (e.g., weighted behavioral cloning) hinders the use of learned value functions, imposing an unnecessary bottleneck on the performance of offline RL.
@@ 5.4 Deep dive 2: Why is DDPG+BC better than AWR?
We have so far seen several empirical results that suggest behavior-regularized policy gradient (e.g., DDPG+BC) should be preferred to weighted behavioral cloning (e.g., AWR) in any cases. What makes DDPG+BC so much better than AWR? There are three potential reasons.
@@
First, AWR only has a mode-covering weighted behavioral cloning term, while DDPG+BC has both mode-seeking first-order value maximization and mode-covering behavioral cloning terms. As a result, actions learned by AWR always lie within the convex hull of dataset actions, whereas DDPG+BC can “hillclimb” the learned value function, even allowing extrapolation to some degree while not deviating too far away from the mode. This not only enables a better use of the value function but yields potentially more optimal actions. To illustrate this, we plot test-time action sampled from policies learned by AWR and DDPG+BC on exorl-walker. Figure 3 shows that AWR actions are relatively centered around the origin, while DDPG+BC actions are more spread out and thus potentially have high optimality.
@@
Takeaway: Current policy extraction can inhibit effective use of the value function Do not use value-weighted behavior cloning (e.g., AWR); always use behavior-constrained policy gradient (e.g., DDPG+BC), regardless of the value learning objective. This enables better scaling of performance with more data and better use of the value function.
@@ 6 Empirical analysis 2: Policy generalization (B3)
We now turn our focus to the third hypothesis, that policy generalization to states that the policy visits at the evaluation time has a significant impact on performance. This is a unique bottleneck to the offline RL problem setting, where the agent encounters new, potentially out-of-distribution states at test time.
@@ 6.1 Analysis setup
One way to measure the degree to which test-time policy generalization affects performance is to evaluate how much room there is for various policy MSE metrics to improve when further training on additional policy rollouts is allowed. The distribution of states induced by rolling out the policy is an ideal distribution to improve performance, as the policy receives direct feedback on its own actions at the states it would visit. Hence, by tracking the extent to which various MSEs improve and how their predictive power towards performance evolves over online interaction, we will be able to understand which is a bigger bottleneck: in-distribution generalization (i.e., improvements towards validation MSE under the offline dataset distribution) or out-of-distribution generalization (i.e., improvements in evaluation MSE under the on-policy state distribution). To this end, we measure these three types of MSEs over the course of online interaction, when learning from a policy trained on offline data only (commonly referred to as the offline-to-online RL setting). Specifically, we train offline-to-online IQL agents on six D4RL [12] tasks (antmaze-{medium, large}, kitchen, and adroit-{pen, hammer, door}), and measure the MSEs with pre-trained expert policies that approximate $\pi^{*}$ (see Section C.3).
@@
6.2 Results: Policy generalization is often the main bottleneck in offline RL
@@ 6.3 Solution 1: Improve offline data coverage
To answer this question, we revert back to our analysis tool of data-scaling matrices from Section 5 and empirically compare the data-scaling matrices on datasets collected by expert policies with different levels of action noises ($\sigma_{\mathrm{data}}$). Figure 7 shows the results of IQL agents on gc-antmaze-large and adroit-pen ($4$ seeds each). The results suggest that the performance of offline RL generally improves as the dataset has better state coverage, despite the increase in suboptimality. This is aligned with our findings in Figure 6, which indicate that the main challenge of offline RL is often not learning an effective policy from suboptimal data, but rather learning a policy that generalizes well to test-time states. In addition, we note that it is crucial to use a value gradient-based policy extraction method (DDPG+BC; see Section 5) in this case as well, where we train a policy from high-coverage data. For instance, in low-data regimes in gc-antmaze-large in Figure 7, AWR fails to fully leverage the value function, whereas DDPG+BC still allows the algorithm to improve performance with better value functions. Based on our findings, we suggest practitioners prioritize high coverage (particularly around the states that the optimal policy will likely visit) over high optimally when collecting datasets for offline RL.
@@ 6.4 Solution 2: Test-time policy improvement
Takeaway: Improving test-time policy accuracy significantly boosts performance Test-time policy generalization is one of the most significant bottlenecks of offline RL. Use high-coverage datasets. Improve policy accuracy on test-time states with on-the-fly policy improvement techniques.
@@ C.2 Data-scaling matrices
We train agents for $1$M steps ($500$K steps for gc-roboverse) with each pair of value learning and policy extraction algorithms. We evaluate the performance of the agent every $100$K steps with $50$ rollouts, and report the performance averaged over the last $3$ evaluations and over $4$ seeds. In Figures 1 and 7, we individually tune the policy extraction hyperparameter ($\alpha$ for AWR and DDPG+BC, and $N$ for SfBC) for each cell, and report the performance with the best hyperparameter. To save computation, we extract multiple policies with different hyperparameters from the same value function (note that this is possible because we use decoupled offline RL algorithms). To generate smaller-sized datasets from the original full dataset, we randomly shuffle trajectories in the original dataset using a fixed random seed, and take the first $K$ trajectories such that smaller datasets are fully contained in larger datasets.
@@ C.6.1 Data-scaling analysis
Default hyperparameters. We mostly follow the original hyperparameters for IQL [24], goal-conditioned IQL [43], and CRL [11]. Tables 1 and 2 list the common and environment-specific hyperparameters, respectively. For SARSA, we use the same implementation as IQL, but with the standard $\ell^{2}$ loss instead of an expectile loss. For pixel-based environments (i.e., gc-roboverse), we use the same architecture and image augmentation as Park et al. [43]. In goal-conditioned environments as well as antmaze tasks, we subtract $1$ from rewards, following previous works [24, 43].
@@ C.6.2 Policy generalization analysis
OPEX $\beta$ $0.3$ (antmaze), $0.0003$ (kitchen)
TTT $\beta$ $0.3$ (antmaze), $5$ (kitchen)

### v2

@@ 1 Introduction
Our main contribution is an analysis of the bottlenecks in offline RL as evaluated via data-scaling properties of various algorithmic choices. Contrary to the conventional belief that value learning is the bottleneck of offline RL algorithms, we find that the performance is often limited by the choice of a policy extraction objective and the degree to which the policy generalizes at test time. This suggests that, with an appropriate policy extraction procedure (e.g., gradient-based policy extraction) and an appropriate recipe for handling generalization (e.g., test-time training with the value function), collecting more high-coverage data to train a value function is a universally better recipe for improving offline RL performance, whenever the practitioner has access to collecting some new data for learning. These results also imply that more research should be pursued in developing policy learning and generalization recipes to translate value learning advances into performant policies.
@@ 3 Main hypothesis
Our primary goal is to understand when and how the performance of offline RL can be bottlenecked in practice. As discussed earlier, there exist three potential factors that could bottleneck an offline RL algorithm: (B1) imperfect value function estimation from data, (B2) imperfect policy extraction from the learned value function, and (B3) imperfect generalization on the test-time states that the policy visits in evaluation rollouts. We note that the bottleneck of an offline RL algorithm under a certain dataset can always be attributed to one or some of these factors, since the policy will attain optimal performance if both value learning and policy extraction are perfect, and perfect generalization to test-time states is possible.
Our main hypothesis in this work is that, somewhat contrary to the prior belief that the accuracy of the value function is the primary factor limiting performance of offline RL methods, policy learning is often the main bottleneck of offline RL. In other words, while value function accuracy is certainly important, how the policy is extracted from the value function (B2) and how well the agent generalizes on states that it visits at the deployment time (B3) are often the main factors that significantly affect both the performance and scalability of offline RL. To verify this hypothesis, we conduct two main analyses in this paper. In Section 5, we compare the effects of value learning and policy extraction on performance under various types of environments, datasets, and algorithms (B1 and B2). In Section 6, we analyze the degree to which the policy generalizes on test-time states affects performance (B3).
@@ 5.2 Results: Policy extraction mechanisms substantially affect data-scaling trends
Table 1: DDPG+BC is often the best policy extraction method. We aggregate the performances over the entire data-scaling matrix and then over $8$ random seeds in each setting. Scores at or above $95\%$ of the best score are highlighted in bold. The table shows that DDPG+BC is better than or as good as AWR in $\mathbf{15}$ out of $\mathbf{16}$ settings. We note that policy extraction hyperparameters are individually tuned for each setting (Figure 12).
Task (Value Algorithm) AWR DDPG+BC SfBC
gc-antmaze-large (IQL) $51$$\pm 2$ $\mathbf{58}$$\pm 2$ $\mathbf{58}$$\pm 1$
gc-antmaze-large (CRL) $37$$\pm 2$ $\mathbf{58}$$\pm 2$ $51$$\pm 2$
antmaze-large (IQL) $12$$\pm 2$ $17$$\pm 4$ $\mathbf{24}$$\pm 3$
antmaze-large (SARSA) $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$ $\mathbf{0}$$\pm 0$
kitchen (IQL) $80$$\pm 1$ $\mathbf{86}$$\pm 1$ $75$$\pm 1$
kitchen (SARSA) $\mathbf{79}$$\pm 1$ $\mathbf{83}$$\pm 1$ $73$$\pm 1$
exorl-walker (IQL) $99$$\pm 1$ $\mathbf{191}$$\pm 6$ $140$$\pm 1$
exorl-walker (SARSA) $94$$\pm 0$ $\mathbf{193}$$\pm 5$ $125$$\pm 1$
Task (Value Algorithm) AWR DDPG+BC SfBC
exorl-cheetah (IQL) $71$$\pm 1$ $\mathbf{101}$$\pm 2$ $77$$\pm 2$
exorl-cheetah (SARSA) $78$$\pm 1$ $\mathbf{131}$$\pm 3$ $89$$\pm 1$
d4rl-hopper (IQL) $\mathbf{53}$$\pm 1$ $\mathbf{52}$$\pm 3$ $43$$\pm 1$
d4rl-hopper (SARSA) $56$$\pm 1$ $\mathbf{61}$$\pm 3$ $50$$\pm 2$
d4rl-walker2d (IQL) $73$$\pm 1$ $\mathbf{81}$$\pm 1$ $68$$\pm 1$
d4rl-walker2d (SARSA) $79$$\pm 0$ $\mathbf{84}$$\pm 0$ $\mathbf{81}$$\pm 1$
gc-roboverse (IQL) $\mathbf{23}$$\pm 2$ $20$$\pm 2$ $14$$\pm 2$
gc-roboverse (CRL) $13$$\pm 1$ $\mathbf{16}$$\pm 2$ $15$$\pm 1$
Figure 1 shows the data-scaling matrices of three policy extraction algorithms (AWR, DDPG+BC, and SfBC) and three value learning algorithms (IQL and {SARSA or CRL}) on eight environments, aggregated from a total of $15{,}488$ runs ($8$ seeds for each cell, numbers after “$\pm$” denote standard deviations). In each matrix, we individually tune the hyperparameter for policy extraction ($\alpha$ or $N$) for each entry. These matrices show how performance varies with different amounts of data for the value and the policy. In our analysis, we specifically focus on the color gradients of these matrices, which reveal the main limiting factor behind the performance of offline RL in each setting. Note that the color gradients are mostly either vertical, horizontal, or diagonal. Vertical ( ) color gradients indicate that the performance is most strongly affected by the amount of policy data, horizontal ( ) gradients indicate it is mostly affected by value data, and diagonal ( ) gradients indicate both.
@@
Among policy extraction algorithms, we find that DDPG+BC almost always achieves the best performance and scaling behaviors across the board, followed by SfBC, and the performance of AWR falls significantly behind the other two extraction algorithms in many cases (Table 1). Notably, the data-scaling matrices of AWR always have vertical ( ) or diagonal ( ) color gradients, implicitly implying that it does not fully utilize the value function (see Section 5.3 for clearer evidence). In other words, a non-careful choice of the policy extraction algorithm (e.g., weighted behavioral cloning) hinders the use of learned value functions, imposing an unnecessary bottleneck on the performance of offline RL.
@@ 5.4 Deep dive 2: Why is DDPG+BC better than AWR?
We have so far seen several empirical results that suggest behavior-regularized policy gradient (e.g., DDPG+BC) should be preferred to weighted behavioral cloning (e.g., AWR) in any case. What makes DDPG+BC so much better than AWR? There are three potential reasons.
@@
First, AWR only has a mode-covering weighted behavioral cloning term, while DDPG+BC has both mode-seeking first-order value maximization and mode-covering behavioral cloning terms. As a result, actions learned by AWR always lie within the convex hull of dataset actions, whereas DDPG+BC can “hillclimb” the learned value function, even allowing extrapolation to some degree while not deviating too far away from the mode. This not only enables a better use of the value function but produces a wider range of actions. To illustrate this, we plot test-time action sampled from policies learned by AWR and DDPG+BC on exorl-walker. Figure 3 shows that AWR actions are relatively centered around the origin, while DDPG+BC actions are more spread out, which can sometimes help achieve an even higher degree of optimality.
@@
Takeaway: Current policy extraction can inhibit effective use of the value function. Do not use value-weighted behavior cloning (e.g., AWR); always use behavior-constrained policy gradient (e.g., DDPG+BC), regardless of the value learning objective. This enables better scaling of performance with more data and better use of the value function.
@@ 6 Empirical analysis 2: Policy generalization (B3)
We now turn our focus to the third hypothesis, that the degree to which the agent generalizes to states that it visits at the evaluation time has a significant impact on performance. This is a unique bottleneck to the offline RL problem setting, where the agent encounters new, potentially out-of-distribution states at test time.
@@ 6.1 Analysis setup
One way to measure the degree to which test-time generalization affects performance is to evaluate how much room there is for various policy MSE metrics to improve when further training on additional policy rollouts is allowed. The distribution of states induced by rolling out the policy is an ideal distribution to improve performance, as the policy receives direct feedback on its own actions at the states it would visit. Hence, by tracking the extent to which various MSEs improve and how their predictive power towards performance evolves over online interaction, we will be able to understand which is a bigger bottleneck: in-distribution generalization (i.e., improvements towards validation MSE under the offline dataset distribution) or out-of-distribution generalization (i.e., improvements in evaluation MSE under the on-policy state distribution). To this end, we measure these three types of MSEs over the course of online interaction, when learning from a policy trained on offline data only (commonly referred to as the offline-to-online RL setting). Specifically, we train offline-to-online IQL agents on six D4RL [12] tasks (antmaze-{medium, large}, kitchen, and adroit-{pen, hammer, door}), and measure the MSEs with pre-trained expert policies that approximate $\pi^{*}$ (see Section C.3).
@@
6.2 Results: Test-time generalization is often the main bottleneck in offline RL
@@ 6.3 Solution 1: Improve offline data coverage
To answer this question, we revert back to our analysis tool of data-scaling matrices from Section 5 and empirically compare the data-scaling matrices on datasets collected by expert policies with different levels of action noises ($\sigma_{\mathrm{data}}$). Figure 7 shows the results of IQL agents on gc-antmaze-large and adroit-pen ($8$ seeds each). The results suggest that the performance of offline RL generally improves as the dataset has better state coverage, despite the increase in suboptimality. This is aligned with our findings in Figure 6, which indicate that the main challenge of offline RL is often not learning an effective policy from suboptimal data, but rather learning a policy that generalizes well to test-time states. In addition, we note that it is crucial to use a value gradient-based policy extraction method (DDPG+BC; see Section 5) in this case as well, where we train a policy from high-coverage data. For instance, in low-data regimes in gc-antmaze-large in Figure 7, AWR fails to fully leverage the value function, whereas DDPG+BC still allows the algorithm to improve performance with better value functions. Based on our findings, we suggest practitioners prioritize high coverage (particularly around the states that the optimal policy will likely visit) over high optimally when collecting datasets for offline RL.
@@ 6.4 Solution 2: Test-time policy improvement
Takeaway: Improving test-time policy accuracy significantly boosts performance. Test-time policy generalization is one of the most significant bottlenecks in offline RL. Use high-coverage datasets. Improve policy accuracy on test-time states with on-the-fly policy improvement techniques.
@@ Appendix A Limitations
Another limitation of our analysis in Section 5 is we only consider policy extraction in continuous-action environments. In discrete-action environments, our takeaway might not directly apply in its current form because (1) DDPG+BC is not straightforwardly defined with discrete actions and (2) it is possible to directly use the Q function to implicitly define a policy (without having a separate policy network). We leave investigating the effect of policy extraction in discrete-action environments for future work.
@@ C.2 Data-scaling matrices
We train agents for $1$M steps ($500$K steps for gc-roboverse) with each pair of value learning and policy extraction algorithms. We evaluate the performance of the agent every $100$K steps with $50$ rollouts, and report the performance averaged over the last $3$ evaluations and over $8$ seeds. In Figures 1 and 7, we individually tune the policy extraction hyperparameter ($\alpha$ for AWR and DDPG+BC, and $N$ for SfBC) for each cell, and report the performance with the best hyperparameter. To save computation, we extract multiple policies with different hyperparameters from the same value function (note that this is possible because we use decoupled offline RL algorithms). To generate smaller-sized datasets from the original full dataset, we randomly shuffle trajectories in the original dataset using a fixed random seed, and take the first $K$ trajectories such that smaller datasets are fully contained in larger datasets.
@@ C.6.1 Data-scaling analysis
Default hyperparameters. We mostly follow the original hyperparameters for IQL [24], goal-conditioned IQL [43], and CRL [11]. Tables 2 and 3 list the common and environment-specific hyperparameters, respectively. For SARSA, we use the same implementation as IQL, but with the standard $\ell^{2}$ loss instead of an expectile loss. For pixel-based environments (i.e., gc-roboverse), we use the same architecture and image augmentation as Park et al. [43]. In goal-conditioned environments as well as antmaze tasks, we subtract $1$ from rewards, following previous works [24, 43].
@@ C.6.2 Policy generalization analysis
OPEX $\beta$ $0.3$ (antmaze), $0.0003$ (kitchen), $0.03$ (d4rl-hopper), $0.1$ (d4rl-walker2d), $1$ (exorl-{walker, cheetah})
TTT $\beta$ $0.3$ (antmaze), $5$ (kitchen), $0.5$ (d4rl-hopper), $0.3$ (d4rl-walker2d), $0.01$ (exorl-{walker, cheetah})